OpenStudy (anonymous):

Find the derivative of H when H(x)= (f(x))^2- (g(x))^2 And f′(x) = 5g(x), g′(x) = 2f(x). A)H'(x)=3(f (x) g(x) B)H'(x)=6(f (x) g(x) C)H'(x)=7(f (x) + g(x) D)H'(x)=7f(x) g(x) E)H'(x)=3(f(x)+g(x)) F)H'(x)=14f(x)g(x)

OpenStudy (john_es):

Use chain rule, \[(H(x))'=(f(x)^2)'-(g(x)^2)'=2f(x)f'(x)-2g(x)g'(x)\] Now, \[f'(x)=5g(x)\\ g'(x)=2f(x)\] You only have to substitute.

OpenStudy (anonymous):

use the chain rule and sub in at the end.

OpenStudy (john_es):

Do you understand it?

OpenStudy (john_es):

After substitute, you'll find, \[H'(x)=6f(x)g(x)\]

OpenStudy (anonymous):

I knew you had to use the chain rule but some times it confuses me lol thank you! @John_ES